The dress store is having a sale where everything is 1/4 off .A woman pays $42 dollars of the merchandise price. What would she have payed at regular price ?
Answer by stanbon(75887)

The dress store is having a sale where everything is 1/4 off .A woman pays $42 dollars of the merchandise price. What would she have payed at regular price ?
-----
Let the regular price be "x":
(3/4)x = 42
x = (4/3)42
x = 4*14 = $56.00 (original price)
==============================
Cheers,
Stan H.